OpenStudy (anonymous):

Help solve this question please: negativeonethird_largem – 7 = 5

OpenStudy (accessdenied):

-(1/3)m - 7 = 5 (we want m alone. lets add 7 to both sides so there's no more 7 next to the m term) -(1/3)m = 12 (then multiply both sides by -3 to cancel out the -(1/3) next to m) m = -36

OpenStudy (campbell_st):

\[-1/3 M - 7 = 5\] add 7 to both sides \[-1/3 M = 12\] multiply both sides by -3 M = -36

OpenStudy (anonymous):

Thank You :)